# In-Context Learning Enhanced Credibility Transformer

### v1

@@ 1.3 Manuscript organization
The next section introduces the ICL-Credibility Transformer. Starting from the base Credibility Transformer, we augment this base architecture by ICL that modifies the instance representations in the latent CLS token space by context batch information. This is achieved by a context token decorator and an ICL transformer. We also discuss the two phase learning procedure, which is common practice in contemporary machine learning tools. Section 3 mainly serves at giving model interpretations and making the connection to Bühlmann credibility. In Section 4, we verify the excellent performance of the ICL-Credibility Transformer on the popular French MTPL data of Dutang et al., (2024). In Section 5, we discuss a zero-shot analysis that shows how our method generalizes to new levels of features that have not been available for training, this reflects a sort of transfer learning using the context. Finally, Section 6 concludes.
@@ 2 In-context learning enhanced Credibility Transformer
This section defines the architecture of the ICL-Credibility Transformer. This architecture is crucially based on the groundbreaking work of Vaswani et al., (2017) on the attention layer mechanism and the Transformer architecture. Based on this seminal work, Richman et al., 2025a introduced the Credibility Transformer which adapts the classical Transformer to tabular data and equips the architecture with a credibility mechanism that enhances learning and prediction. We will use the notation of Richman et al., 2025a .
@@ 2.1 ICL-Credibility Transformer architecture
    an outcome token decorator that encodes claims information;
@@
    a (frozen) decoder that produces the final predictions.
@@ 2.1.1 Credibility Transformer encoder with CLS token
Our base Credibility Transformer (CT) follows the architecture introduced by Richman et al., 2025a , and we also use the notation of that reference. It takes an input $\boldsymbol{x}$ and it outputs a prediction $\mu^{\rm CT}(\boldsymbol{x})$. For this, it encodes the input $\boldsymbol{x}$ into a learnable credibilitized CLS token
@@
for some encoding dimension $2b\in{\mathbb{N}}$; see Richman et al., 2025a (, Formula (2.12)). This credibilitized CLS token is a $2b$-dimensional encoding of the feature $\boldsymbol{x}$ (containing positional encodings). It is obtained by a Transformer architecture which is extended by the CLS token of Devlin et al. Devlin et al., (2019). During model fitting this CLS token is obtained by a credibility weighted average of individual instance information and global (prior) population information; see Richman et al., 2025a (, Section 2.2).
The second part of the Credibility Transformer is the decoder $\boldsymbol{z}^{\rm decod}$ which maps the credibilitized CLS token to the prediction. Composing this encoder and decoder pair provides the base Credibility Transformer
@@
This is precisely the predictive model considered in Richman et al., 2025a (, Section 2.2), and in Table 2 of that reference it is verified on the popular French MTPL insurance dataset that this architecture provides an excellent predictive performance.
Broadly speaking, for all following steps, we keep the decoder layer (2.2) of the trained base Credibility Transformers fixed, i.e., we freeze the estimated weights of the fitted decoder $\widehat{\boldsymbol{z}}^{\rm decod}$, but before we apply this frozen decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ on the credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}=\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x})$, we enrich these tokens by in-context information. We will start the ICL training from the weights of the encoder part of the model but we allow these to be optimized to produce a CLS token that is optimal for ICL.
@@ 2.1.2 Outcome token decorator
In the next step, we insert a module into the pre-calibrated base Credibility Transformer (2.2) that enables ICL. For this additional step, we assume that there two kinds of data available. We have a context batch and we have a target batch. The target batch collects the instances whose responses need to be predicted, and this prediction is supported not only by the features of the target batch, but also by context data for which the responses are available.
We denote the context batch by ${\cal B}_{\rm context}=(Y_{j},\boldsymbol{x}_{j},v_{j})_{j\in{\cal I}_{\rm context}}$ and the target batch is denoted by ${\cal B}_{\rm target}=(Y_{i},\boldsymbol{x}_{i},v_{i})_{i\in{\cal I}_{\rm target}}$. These two sets are disjoint to avoid any leakage of information. The outcome token decorator is used to enrich the credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{j})$ of the context set $j\in{\cal I}_{\rm context}$ by outcome information $Y_{j}$ (which we assume is available only on the context batch). This will then serve as the context to predict the instances in the target batch ${\cal B}_{\rm target}$, where we assume that only the feature information $\boldsymbol{x}_{i}$ and its encoded version $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{i})$ is available to predict the response $Y_{i}$.
To simplify the notation, we merge the two index sets ${\cal I}={\cal I}_{\rm target}\cup{\cal I}_{\rm context}$ assuming that these indices allow for a unique identification of all instances. We then introduce the mask for $i\in{\cal I}$
@@
${\cal B}=(Y_{i},\boldsymbol{x}_{i},v_{i},M_{i})_{i\in{\cal I}}.$
@@
The decorated tokens are then defined by
@@
  • credibilitized tokens $\widehat{\boldsymbol{c}}^{\rm cred}$, which, as above, have been produced using the encoder part of the original Credibility Transformer;
  • we decorate all instances $i\in{\cal I}_{\rm context}$ from the context batch ${\cal B}_{\rm context}$ with encoded target information $\boldsymbol{z}^{\rm FNN1}(Y_{i})$;
  • we integrate this target information using Bühlmann, (1967) credibility, assigning a higher credibility for bigger case weights $v_{i}>0$, and the credibility coefficient $\kappa>0$ regulates the level of information considered; $\kappa$ is a hyper-parameter that is either chosen by the modeler or learned during training.
@@ 2.1.3 ICL Transformer
we also refer to Figure 1. The goal is to predict all the responses $Y_{i}$ of the target set $i\in{\cal I}_{\rm target}$ in one shot using the entire context batch ${\cal B}_{\rm context}$. To ensure independence within the target set in the next step, we need to modify the previously introduced mask to
@@
This masks all pairs $(i,j)$ in the target batch with -$\infty$ if their indices $i\neq j$ differ. This will imply that the instances in the target batch cannot interact. We define the mask tensor $\mathbf{M}=(M^{\infty}_{i,j})_{i,j\in{\cal I}}$, and we set $n=|{\cal I}|$ for the total batch size and $n_{1}=|{\cal I}_{\rm target}|$ for the target batch size.
Next, we define the causal ICL Transformer layer. We start by defining a causal attention layer; see Vaswani et al., (2017). Using three different time-distributed FNNs, we generate the query, key and value tensors, respectively,
@@
$\mathbf{A}=\mathbf{A}(\mathbf{Q},\mathbf{K},\mathbf{M})=\operatorname{softmax}\left(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{2b}+\mathbf{M}\right)~\in~{\mathbb{R}}^{n\times n},$
@@
The ICL Transformer layer includes two skip connections, another time-distributed FNN $\boldsymbol{z}^{\rm t-FNN2}$, and dropout layers and layer-normalization layers. The first regularized skip connection is given by
@@
This then allows one to compute the ICL Transformer output
$\mathbf{C}^{\rm ICL-trans}=\operatorname{LayerNorm}\left(\mathbf{C}^{\rm mid}+\operatorname{Dropout}\left(\boldsymbol{z}^{\rm t-FNN2}\left(\mathbf{C}^{\rm mid}\right)\right)\right)~\in~{\mathbb{R}}^{n\times 2b}.$ (2.6)
@@ 2.1.4 Frozen decoder
The last step is to decode the ICL Transformer tensor $\mathbf{C}^{\rm ICL-trans}$ to form the predictions on the target batch, i.e., for the responses $Y_{i}$, $i\in{\cal I}_{\rm target}$. For this we use the frozen decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ from the first training phase. This architectural constraint ensures that all ICL computations operate within the original representation space learned during the first learning phase, which provides some theoretical elegance and practical benefits that maintain the interpretability of the original space in which the CLS tokens live.
Let us denote the row vectors of the tensor $\mathbf{C}^{\rm ICL-trans}$ as follows
$\mathbf{C}^{\rm ICL-trans}=\left[\boldsymbol{c}^{\rm ICL-trans}_{1},\ldots,\boldsymbol{c}^{\rm ICL-trans}_{n_{1}},\boldsymbol{c}^{\rm ICL-trans}_{n_{1}+1},\ldots,\boldsymbol{c}^{\rm ICL-trans}_{n}\right]^{\top}~\in~{\mathbb{R}}^{n\times 2b},$
@@
$\widehat{\mu}^{\rm ICL-CT}(\boldsymbol{x}_{i};{\cal B}_{\rm context})=\widehat{\boldsymbol{z}}^{\rm decod}\left(\boldsymbol{c}^{\rm ICL-trans}_{i}\right),\qquad\text{ for $i\in{\cal I}_{\rm target}$}.$ (2.7)
We give some remarks.
@@
  • Implicit regularization: By constraining the ICL mechanism to operate within the learned representation space, we prevent it from arbitrarily transforming representations in ways that might over-fit to the ICL training objective. This acts as a strong regularization mechanism.
@@
  • Information bottleneck: From a more information-theoretic perspective, the frozen decoder enforces an information bottleneck that prevents the ICL mechanism from encoding arbitrary information that was not present in the original representation space. This aligns with the principle of minimal sufficient statistics, ensuring that the ICL mechanism focuses on extracting and utilizing the most relevant contextual information for improving predictions. On the other hand, this also requires that the learned credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}$ are sufficiently rich, otherwise the ICL fine-tuning will not have any positive effect on improving the Credibility Transformer predictions.
@@
    the encoder of the original Credibility Transformer;
@@
${\cal L}_{\text{ICL}}\left({\cal B}_{\rm target};{\cal B}_{\rm context}\right)=\frac{1}{|{\cal I}_{\rm target}|}\sum_{i\in{\cal I}_{\rm target}}v_{i}\,L\left(Y_{i},\widehat{\mu}^{\rm ICL-CT}(\boldsymbol{x}_{i};{\cal B}_{\rm context})\right),$ (2.8)
for a suitable loss function $L$ (comparing responses $Y$ and predictions $\widehat{\mu}$) and case weights $v>0$. In practice, we will have many target-context pairs, and this second training phase performs gradient descent fine-tuning on all these pairs, averaging the losses (2.8) of the the available target-context pairs.
@@ 2.2 Learning procedure
The fitting procedure presented above has two different learning phases. The first one is used to learn the credibilitized CLS token encoder $\widehat{\boldsymbol{c}}^{\rm cred}$ and the decoder $\widehat{\boldsymbol{z}}^{\rm decod}$, respectively, and the second phase is used to learn the context augmentation involving the context token decorator and the ICL Transformer, as well as fine-tuning the CLS token decorator. This two step approach is employed for computational reasons and for robustifying the training procedure. There is one critical point in this two stage procedure though, namely, one needs to ensure that the learned credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{i})$, $i\in{\cal I}$, are sufficiently rich and carry sufficient information of the inputs $\boldsymbol{x}_{i}$; an encoding can be thought of a data compression. The critical point is that the training of the encoder $\widehat{\boldsymbol{c}}$ and the decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ only focuses on optimal predictive performance, without accounting for the space of encoded features $(\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}))_{i\in{\cal I}}$ being still sufficiently rich to be able to benefit from ICL in the second step. Basically, there are two ways of ensuring this. First, certainly the embedding dimension $2b$ of the CLS tokens should be sufficiently large, otherwise we are facing the bottleneck already in the very beginning. Second, one can employ a third training phase, allowing for fine-tuning of all involved weights by jointly training the entire architecture starting from the pre-trained weights of the first two phases. To ensure stability of the learned representation space, this third (fine-tuning) step should be constrained by a strong regularization or it should only consider very few gradient descent steps with moderate step sizes.
@@
Clearly, to benefit from context data, the target batch and the context batch have to be associated. This association can reflect target-context pairs that consider the same products, the same local regions or the same circumstances in which these instances operate, without having these (soft) factors encoded in the features. Such examples are the most obvious ones in which the ICL-Credibility Transformer can successfully operate. A less obvious though similar set-up is the problem of having a high-cardinality categorical feature, e.g., vehicle models within a motor insurance portfolio. Categorical encoding (one-hot encoding, dummy coding or entity embedding) will then involve a vast number of parameters which cannot be estimated reliably (with credibility). Instead of trying to estimate a multitude of parameters, we rather could use similar instances as the context batch. Either there is a natural association or one can try to construct a neighborhood of similar instances using an unsupervised learning method; see Wüthrich et al., (2025, Chapter 10).

### v2

@@ 1.3 Manuscript organization
The next section introduces the ICL-Credibility Transformer. Starting from the base Credibility Transformer, we augment this base architecture by ICL that modifies the instance representations in the latent CLS token space by context batch information. This is achieved by a context token decorator and an ICL transformer. We also discuss the two phase learning procedure, which is common practice in contemporary machine learning tools. Section 3 mainly serves at giving model interpretations and making the connection to Bühlmann credibility. In Section 4, we verify the excellent performance of the ICL-Credibility Transformer on the popular French MTPL data of Dutang et al. (2024). In Section 5, we discuss a zero-shot analysis that shows how our method generalizes to new levels of features that have not been available for training. Finally, Section 6 concludes.
@@ 2 In-context learning enhanced Credibility Transformer
This section defines the architecture of the ICL-Credibility Transformer. This architecture is crucially based on the groundbreaking work of Vaswani et al. (2017) on the attention layer mechanism and the Transformer architecture. Based on this seminal work, Richman et al. (2025a) introduced the Credibility Transformer which adapts the classical Transformer to tabular data and equips the architecture with a credibility mechanism that enhances learning and prediction. This outlline is based on the notation of Richman et al. (2025a).
@@ 2.1 ICL-Credibility Transformer architecture
    an outcome token decorator that encodes claims information of the context batch;
@@
    a (frozen) decoder that produces the final predictions from the tokens.
@@ 2.1.1 Credibility Transformer encoder with CLS token
Our base Credibility Transformer (CT) follows the architecture introduced by Richman et al. (2025a), and we also use the notation of that reference. It takes an input $\boldsymbol{x}$ and it outputs a prediction $\mu^{\rm CT}(\boldsymbol{x})$ by an encoder-decoder architecture. The encoder $\boldsymbol{c}^{\rm cred}$ takes an input $\boldsymbol{x}$ and maps it to the credibilitized CLS token
@@
for a fixed encoding dimension $2b\in{\mathbb{N}}$; see Richman et al. (2025a, Formula (2.12)). This credibilitized CLS token is a $2b$-dimensional encoding of the feature $\boldsymbol{x}$ (containing positional encodings). It is obtained by a Transformer architecture which is extended by the CLS token of Devlin et al. Devlin et al. (2019). During model fitting this CLS token is computed by a credibility weighted average of individual instance information and global (prior) population information; for full technical details we refer to Richman et al. (2025a, Section 2.2).
The second part of the Credibility Transformer is the decoder $\boldsymbol{z}^{\rm decod}$ which maps the credibilitized CLS token to the prediction, usually we take a FNN architecture for the decoding step. Composing this encoder-decoder pair provides the base Credibility Transformer
@@
This base Credibility Transformer architecture is described in full detail in Richman et al. (2025a, Section 2).
@@
This is precisely the predictive model considered in Richman et al. (2025a, Section 2), and in Table 2 of that reference verifies the excellent predictive performance of this architecture on the popular French MTPL insurance dataset.
Broadly speaking, for the following steps, we keep the decoder architecture $\widehat{\boldsymbol{z}}^{\rm decod}$ of the trained base Credibility Transformer (2.2) fixed, i.e., we freeze the estimated weights of the fitted decoder $\widehat{\boldsymbol{z}}^{\rm decod}$, and before applying this frozen decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ on the credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}=\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x})$, we enrich these by in-context information.
@@ 2.1.2 Outcome token decorator
In the next step, we insert a module into the pre-calibrated base Credibility Transformer (2.2) that enables ICL. For this additional step, we assume that there two kinds of data available, namely, we assume having a context batch and a target batch. The target batch collects the instances whose responses need to be predicted. This prediction is supported not only by the features of the target batch, but also by context data for which the responses are available.
We denote the context batch by ${\cal B}_{\rm context}=(Y_{j},\boldsymbol{x}_{j},v_{j})_{j\in{\cal I}_{\rm context}}$, and the target batch is denoted by ${\cal B}_{\rm target}=(Y_{i},\boldsymbol{x}_{i},v_{i})_{i\in{\cal I}_{\rm target}}$. These two sets are disjoint to avoid any leakage of information. The outcome token decorator is used to enrich the credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{j})$ of the context set $j\in{\cal I}_{\rm context}$ by outcome (response) information $Y_{j}$, which is assumed to be available only on the context batch. This will then serve as the context to predict the responses in the target batch ${\cal B}_{\rm target}$.
To simplify the notation, we merge the two index sets ${\cal I}={\cal I}_{\rm target}\cup{\cal I}_{\rm context}$ assuming that these indices allow for a unique identification of all instances. We then introduce the mask
@@
${\cal B}=(Y_{i},\boldsymbol{x}_{i},v_{i},M_{i})_{i\in{\cal I}},$
with response $Y_{i}$ only being available for $M_{i}=1$.
@@
The decorated token is then defined by
@@
  • credibilitized CLS token $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{i})$, which, as above, has been produced using the encoder part of the base Credibility Transformer;
  • we decorate all instances $i\in{\cal I}_{\rm context}$ from the context batch ${\cal B}_{\rm context}$, $M_{i}=1$, with encoded target information $\boldsymbol{z}^{\rm FNN1}(Y_{i})$;
  • we integrate this target information using Bühlmann (1967) credibility, assigning a higher credibility for bigger case weights $v_{i}>0$, and the credibility coefficient $\kappa>0$ regulates the level of information considered; $\kappa$ is a hyper-parameter that is either chosen by the modeler or optimized by grid search.
@@ 2.1.3 ICL Transformer
we also refer to Figure 1. The goal is to predict all responses $Y_{i}$ of the target set $i\in{\cal I}_{\rm target}$ in one shot using the entire context batch ${\cal B}_{\rm context}$. To ensure independence within the target set in the next step, we introduce the mask operator (on the log-scale)
@@
This masks all pairs $(i,j)$ in the target batch ${\cal B}_{\rm target}$ with -$\infty$ if their indices $i\neq j$ differ. This will imply that the instances in the target batch cannot interact. We define the mask tensor $\mathbf{M}=(M^{\infty}_{i,j})_{i,j\in{\cal I}}$, and we set $n=|{\cal I}|$ for the total batch size and $n_{1}=|{\cal I}_{\rm target}|$ for the target batch size.
Next, we define the causal ICL Transformer layer. We start by defining a causal attention layer; see Vaswani et al. (2017). Using three different time-distributed FNNs, we generate the query, key and value tensors, respectively,
@@
$\mathbf{A}=\mathbf{A}(\mathbf{Q},\mathbf{K},\mathbf{M})=\operatorname{softmax}\left(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{2b}+\mathbf{M}\right)~\in~{\mathbb{R}}^{n\times n},$ (2.6)
@@
The ICL Transformer layer includes two skip connections, another time-distributed FNN $\boldsymbol{z}^{\rm t\text{-}FNN2}$, and dropout layers and layer-normalization layers. The first regularized skip connection is given by
@@
This then allows one to compute the ICL Transformer tensor
$\mathbf{C}^{\rm ICL\text{-}trans}=\operatorname{LayerNorm}\left(\mathbf{C}^{\rm mid}+\operatorname{Dropout}\left(\boldsymbol{z}^{\rm t\text{-}FNN2}\left(\mathbf{C}^{\rm mid}\right)\right)\right)~\in~{\mathbb{R}}^{n\times 2b}.$ (2.7)
Remark that (2.7) is the classical Transformer encoder introduced by Vaswani et al. (2017), and the mask operator $\mathbf{M}$ in (2.6) and the time-distributed FNN layers ensure that different instances of the target batch do not interact in this ICL Transformer encoder.
@@ 2.1.4 Frozen decoder
The last step is to decode the ICL Transformer tensor $\mathbf{C}^{\rm ICL\text{-}trans}$ to form the predictions on the target batch, i.e., for the responses $Y_{i}$, $i\in{\cal I}_{\rm target}$. For this we use the frozen decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ from the first training phase. This architectural constraint ensures that all ICL computations operate within the original representation space learned during the first learning phase. This provides the practical benefit of maintaining the interpretability of the original CLS token space.
Let us denote the row vectors of the ICL Transformer tensor $\mathbf{C}^{\rm ICL\text{-}trans}$ as follows
$\mathbf{C}^{\rm ICL\text{-}trans}=\left[\boldsymbol{c}^{\rm ICL\text{-}trans}_{1},\ldots,\boldsymbol{c}^{\rm ICL\text{-}trans}_{n_{1}},\,\boldsymbol{c}^{\rm ICL\text{-}trans}_{n_{1}+1},\ldots,\boldsymbol{c}^{\rm ICL\text{-}trans}_{n}\right]^{\top}~\in~{\mathbb{R}}^{n\times 2b},$
@@
$\widehat{\mu}^{\rm ICL\text{-}CT}(\boldsymbol{x}_{i};{\cal B}_{\rm context})=\widehat{\boldsymbol{z}}^{\rm decod}\left(\boldsymbol{c}^{\rm ICL\text{-}trans}_{i}\right),\qquad\text{ for $i\in{\cal I}_{\rm target}$}.$ (2.8)
We give remarks that all relate to the frozen decoder.
@@
  • Implicit regularization: By constraining the ICL mechanism to operate within the learned representation space, we prevent it from arbitrarily transforming representations in ways that might over-fit to the ICL training objective, i.e., this acts as a regularization mechanism. This can also be seen as an information bottleneck that prevents the ICL mechanism from encoding arbitrary context information.
@@
    the encoder of the base Credibility Transformer;
@@
${\cal L}_{\text{ICL}}\left({\cal B}_{\rm target};{\cal B}_{\rm context}\right)=\frac{1}{|{\cal I}_{\rm target}|}\sum_{i\in{\cal I}_{\rm target}}v_{i}\,L\left(Y_{i},\widehat{\mu}^{\rm ICL\text{-}CT}(\boldsymbol{x}_{i};{\cal B}_{\rm context})\right),$ (2.9)
for a suitable loss function $L$ (comparing the responses $Y$ and predictions $\widehat{\mu}$) and case weights $v>0$. In practice, we will have many target-context pairs, and this second training phase performs gradient descent fine-tuning on all these pairs, averaging the losses (2.9) of the available target-context pairs.
@@ 2.2 Learning procedure
The fitting procedure presented above has two different learning phases. Phase 1: The first phase is used to learn the credibilitized CLS token encoder $\widehat{\boldsymbol{c}}^{\rm cred}$ and the decoder $\widehat{\boldsymbol{z}}^{\rm decod}$, respectively. Phase 2: The second phase is used to learn the context augmentation involving the context token decorator and the ICL Transformer, as well as fine-tuning the CLS token decorator. This two step approach is employed for robustifying the training procedure, but it also has computational advantages. There is one critical point in this two stage procedure though, namely, one needs to ensure that the learned credibilitized CLS tokens $\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}_{i})$, $i\in{\cal I}$, are sufficiently rich and carry sufficient information of the inputs $\boldsymbol{x}_{i}$; an encoding can be thought of a data compression. The critical point is that the training of the encoder $\widehat{\boldsymbol{c}}^{\rm cred}$ and the decoder $\widehat{\boldsymbol{z}}^{\rm decod}$ only focuses on optimal predictive performance, without accounting for the space of encoded features $(\widehat{\boldsymbol{c}}^{\rm cred}(\boldsymbol{x}))_{i\in{\cal I}}$ being still sufficiently rich to be able to benefit from the ICL mechanism in the second step. Basically, there are two ways of ensuring this. First, certainly the embedding dimension $2b$ of the CLS tokens should be sufficiently large, otherwise we are facing the bottleneck already in the very beginning. Second, one can employ a third training phase. This motivates Phase 3: This third phase allows for a fine-tuning of all involved weights by jointly training the entire architecture starting from the pre-trained weights of the first two phases. To ensure stability of the learned representation space, this third (fine-tuning) step should be constrained by a strong regularization or it should only consider very few gradient descent steps with moderate step sizes (learning rates).
@@
Clearly, to benefit from context data, the target batch and the context batch have to be associated. This association can reflect target-context pairs that consider the same products, the same local regions or the same circumstances in which these instances operate, without having these (soft) factors encoded in the features. Such examples are the most obvious ones in which the ICL-Credibility Transformer can successfully operate. A less obvious though similar set-up is the problem of having a high-cardinality categorical feature, e.g., vehicle models within a motor insurance portfolio. Categorical encoding (one-hot encoding, dummy coding or entity embedding) will then involve a vast number of parameters which cannot be estimated reliably (with credibility). Instead of trying to estimate a multitude of parameters, we rather could use similar instances as the context batch. Either there is a natural association or one can try to construct a neighborhood of similar instances using an unsupervised learning method; see Wüthrich et al. (2025, Chapter 10).